OpenStudy (anonymous):

How many solutions does this system have? 2x+y=3 6x=9-3y a.1 b. none c.infinite d.2

OpenStudy (hoblos):

2x+y=3 6x=9-3y first equation --> y = -2x+3 second equation ----> 3y = -6x+9 y = -2x +3 the two equations are the same what would be the solution for a system of identical equations ?

OpenStudy (anonymous):

im thinking c?

OpenStudy (hoblos):

correct it's c

OpenStudy (anonymous):

which cannot describe a system of linear equations? a. no solution b. exactly 2solutions c. infinite solutions d. exactly 1 solution

OpenStudy (hoblos):

linear equations represent straight lines do you know what is the intersection of two straight lines ?

OpenStudy (anonymous):

two solution?

OpenStudy (hoblos):



OpenStudy (hoblos):

how many points of intersection ?

OpenStudy (anonymous):

1

OpenStudy (hoblos):

yeah so there is one solution for a system of linear equations

OpenStudy (anonymous):

so d

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

it was exactly two solutions for people doing Lesson 1 Solving Systems by Graphing, the answers are: 1.D 2.C 3.D 4.B I got a 75%, that should help. :/